# Quantifying Classifier Utility under Local Differential Privacy

### v2

@@ 5.1 Robustness Hyperrectangle
The commonly used robustness radius for classifiers [18] is a conservative measure of robustness. It defines the robustness area as an $\ell_{\infty}$-ball, [^6] Other $\ell_{p}$-balls, such as $\ell_{1}$-ball and $\ell_{2}$-ball, are also used, but they are also symmetric across all dimensions. where all dimensions share the same radius. However, in practice, different dimensions often have varied robustness radii, making the true robustness area larger than that defined by an $\ell_{\infty}$-ball. In our context, accurately describing the robustness area is crucial for refining utility quantification. To this end, we define the robustness hyperrectangle, which accounts for the robustness of each dimension individually. The formal definition is as follows.
@@
The robustness radius is a special case of the robustness hyperrectangle, where each dimension has the same size $\theta$ around $x$. More generally, we say that the classifier $h$ is $\theta_{\diamond}$-locally robust at $x$.
Finding robustness hyperrectangle. The robustness hyperrectangle $\theta_{\diamond}$ is not unique, unlike the robustness radius. Different methods for finding $\theta_{\diamond}$ may yield different results. Heuristically, we aim to include the robustness radius to ensure an improved utility quantification.
@@
Figure 5: Two robustness hyperrectangles (green dashed boxes) for the classifier in Figure 3. The original robustness radius $\theta=0.2$ corresponds to the gray dashed box.
@@
Each dimension of $\theta_{\diamond}$, represented as $a_{i},b_{i}]$, influences the concentration analysis of LDP mechanisms. Larger $a_{i},b_{i}]$ allows for greater perturbation, leading to a stronger utility guarantee. Specifically, the concentration level $\theta$ in the utility guarantee $\rho(\varepsilon,\theta)$ is replaced with the refined $\theta_{\diamond}$, resulting in the improved utility guarantee $\rho(\varepsilon,\theta_{\diamond})$.
Refined utility quantification. Given a $d$-dimensional classifier $h$, LDP mechanism $\mathcal{M}_{\varepsilon}$ for each dimension, and private data $x$. Denote the utility quantification $\rho(\varepsilon,\theta)$ as defined in Section 4.3. It can be refined to: With probability at least $\rho(\varepsilon,\theta_{\diamond})$, the classifier $h$ preserves the correct classification result under the input perturbed by $d$ independent $\mathcal{M}_{\varepsilon}$ (i.e. pure $d\varepsilon$-LDP), where [^7] We omit the term $(1-\omega)(1-\tau)$ in $\rho(\varepsilon,\theta_{\diamond})$ for simplicity in presentation. It is almost $1$ in practice (e.g. $\omega=0.05,\tau=0.01$) and thus also negligible.
@@
5.2 PAC Privacy
Probably approximately correct (PAC) privacy [62] provides a formal framework for relaxing the worst-case privacy guarantee. While pure LDP requires the privacy constraint to hold in all cases, PAC privacy allows a small probability of failure, i.e. probabilistic LDP. This relaxation allows the use of a wider range of mechanisms, notably the Gaussian mechanism, which is widely used in privacy-preserving machine learning.
To simplify notation, we denote the privacy loss [17] between $x_{1}$ and $x_{2}$ at $\tilde{x}$ as
@@ 5.2 PAC Privacy
Pure $\varepsilon$-LDP can then be expressed as: $\mathcal{L}_{\mathcal{M},x_{1},x_{2}}(\tilde{x})\leq\varepsilon$ for all $x_{1},x_{2},\tilde{x}$. In contrast, $(\varepsilon,\delta)$-PAC LDP is defined by a relaxed $\delta$-probabilistic condition. [^8] Readers may wonder about the relationship between $(\varepsilon,\delta)$-PAC LDP and $(\varepsilon,\delta)$-LDP. In fact, the traditional $(\varepsilon,\delta)$-LDP from Dwork [17] is ambiguous and has many interpretations. Concentrated privacy [15] from Dwork, Renyi privacy [46], and PAC privacy [62] can address this ambiguity. See Appendix C.3 for a detailed discussion.
Definition 10 (PAC LDP, adopted from [62]).
@@
Combination of PAC LDP mechanisms. When applying $d$ independent PAC LDP mechanisms to $d$ dimensions, if we follow the combination theorem from Dwork [17], it gives $(d\varepsilon,d\delta)$-PAC LDP. However, this result is not tight, as it follows the failure probability $d\delta$ can exceed $1$, which is impossible. We provide a tighter result for combining $d$ independent PAC LDP mechanisms.
@@ 5.2.1 Privacy Indicator
Refined utility quantification. Given a $d$-dimensional classifier $h$, combined LDP mechanism $\mathcal{M}^{d}_{\varepsilon}$, privacy indicator $\mathcal{I}_{\delta}(\cdot)$, and raw data $x$, the utility quantification $\rho(\varepsilon,\theta)$ can be refined to: With probability at least $\delta+(1-\delta)\rho(\varepsilon,\theta_{\diamond})$, the classifier $h$ preserves the correct result under the input perturbed by $\mathcal{I}_{\delta}(\mathcal{M}^{d}(x))$, which guarantees $(d\varepsilon,\delta)$-PAC LDP.
@@
5.2.2 Extended Gaussian Mechanism
It is already known that the Gaussian mechanism [17] cannot satisfy pure $\varepsilon$-LDP, making the privacy indicator inapplicable. Nonetheless, Dwork has shown that it can satisfy $(\varepsilon,\delta)$-LDP for $\varepsilon\leq 1$. Under the notion of PAC LDP, we provide an extended Gaussian mechanism to work for any $\varepsilon$. [^9] Analytical Gaussian mechanism [5] works for any $\varepsilon$, but it is based on an alternative privacy definition and lacks analytical form of the noise scale $\sigma$. Appendix B.2.3 provides the detailed comparison. The following theorem is the form for data domain $x\in[0,1]$.
@@ 5.2.2 Extended Gaussian Mechanism
(Sketch) The proof generally follows the structure of Dwork et al. [17], but uses a rewritten privacy loss and a different tail bound for the Gaussian distribution. Appendix B.2 provides the details and compares this bound with others. ∎
@@
Refined utility quantification. Given a classifier $h$, combined Gaussian mechanism $\mathcal{M}^{d}$ with $\delta$, and $d$-dimensional private data $x$, the utility quantification $\rho(\varepsilon,\theta_{\diamond})$ can be refined to: With probability at least $\rho(\varepsilon,\theta_{\diamond})$, the classifier $h$ preserves the correct classification result under the input perturbed by $\mathcal{M}^{d}$, which guarantees $(d\varepsilon,1-(1-\delta)^{d})$-PAC LDP.
The extended Gaussian mechanism and the privacy indicator provide two approaches to achieve $(\varepsilon,\delta)$-PAC LDP. While both operate under the same privacy notion, the Gaussian mechanism actually provides a stronger privacy: when it fails to satisfy the pure $\varepsilon$-LDP (with probability $\delta$), the privacy loss exceeds $\varepsilon$ but remains bounded by a larger value. In contrast, the privacy indicator exposes the original data when it fails.
6 Discussions
This section discusses detailed privacy questions in the utility quantification framework and the complexity of finding the robustness radius.
@@ 6 Discussions
In the LDP model, it is assumed that the classifier is curious about private data but honest in reporting results, i.e. it outputs $h(\mathcal{M}(x))$ rather than fabricating results. This is a weak assumption, as any utility evaluation becomes meaningless if the classifier fabricates results.
Utility quantification, as a different scenario from using the classifier, focuses on a specific input $x$ but does not involve interaction with users’ data. It can be conducted either by the classifier designer or by the users, and in both cases, there is no privacy leakage for the users: (i) If conducted by the classifier designer, e.g. for improving design, there is no interaction with users, and thus no privacy concerns arise. (ii) If conducted by the users, e.g. to achieve a better privacy-utility trade-off for their data, they can obtain the robustness radii without revealing their original data. For white-box classifiers, users know the classifier parameters and can compute the robustness radii. For black-box classifiers, users can uniformly query the classifier over the entire input domain (i.e. prediction for nearly all input $x\in[0,1]^{d}$), without revealing their specific data. This one-time procedure yields robustness radii for all inputs and privacy parameters. Thus, the robustness radius at any input can be considered public knowledge, which is a reasonable assumption.
Complexity of Finding the Robustness Radius
7 Case Studies
@@ 7.1 Setup
  • Stroke Prediction [19]: Predicts stroke likelihood based on features like age, hypertension, and BMI. This dataset contains $6$ numerical input dimensions and a binary output. The sensitive features are “Age” and “BMI”.
  • Bank Customer Attrition [42]: Predicts whether a bank customer will leave based on features like age, salary, and tenure. This dataset has $9$ numerical input dimensions and a binary output. The sensitive features are “Age” and “Estimated Salary”.
  • MNIST-$\times7$ (variant of [12]): Predicts handwritten digits (0 to $9$) based on pixel values. This dataset has $49$ input dimensions and a $10$-class output. All dimensions are treated as sensitive features.
The first two datasets involve direct sensitive user inputs, while the third is a standard image classification dataset. For the first two datasets, we use the sklearn library with cross-validation to train Logistic Regression and Random Forest classifiers. These traditional classifiers are widely used in medical and financial domains due to their interpretability and lower data requirements [2]. For the third dataset, which consists of high-dimensional images, we use the torch library to train a Convolutional Neural Network (CNN) classifier. All datasets are normalized to $[0,1]$ for training, as normalization aids classifier convergence and is a common practice.
@@
Table 1: Summary of LDP mechanisms, types of classifiers, and datasets used in the case studies.
Mechanism
Laplace [17], Gaussian (Theorem 4),
PM [54], SW [37], Exponential [48], $k$-$R$ [31]
Classifier
Low-dim: Logistic Regression, Random Forest;
High-dim: Neural Network
Dataset
Low-dim: Stroke Prediction [19], Bank Customer
Attrition [42]; High-dim: MNIST-$\times7$ (variant of [12])
We present theoretical utility quantification for the classifiers under the LDP mechanisms discussed above. This quantification is based on the robustness hyperrectangle $\theta_{\diamond}$ of the classifier at a randomly selected record, determined using the search method described in Section 5.1. We also compare the theoretical results with empirical utility observations. Specifically, for a given classifier and mechanism, we provide and compare:
@@
7.2 Utility Quantification Results
7.2.1 Stroke Prediction
@@
$\begin{aligned} \rho(\varepsilon,\theta)=&[F_{\text{PM}_{\varepsilon,\text{Age}}}(1)-F_{\text{PM}_{\varepsilon,\text{Age}}}(0.63)]\cdot[F_{\text{PM}_{\varepsilon,\text{BMI}}}(1)-F_{\text{PM}_{\varepsilon,\text{BMI}}}(0)]\\ =&1-F_{\text{PM}_{\varepsilon,\text{Age}}}(0.63).\end{aligned}$
@@
Figure 6: Empirical and theoretical utility for two classifiers trained on the Stroke Prediction dataset.
@@
The Random Forest classifier exhibits a larger gap between $\rho(\varepsilon,\theta_{\diamond})$ and $\hat{\rho}(\varepsilon)$ compared to the Logistic Regression classifier. This discrepancy arises from the robustness hyperrectangle $\theta_{\diamond}$, which is a conservative estimate of the overall robustness region (difficult to determine precisely). For Logistic Regression, the decision boundary is a hyperplane [50], making it more predictable and easier to approximate a hyperrectangle, resulting in a more accurate $\rho(\varepsilon,\theta_{\diamond})$. Appendix C.4.3 provides additional details on decision boundaries.
@@
Table 2: Time cost comparison (in milliseconds).
@@
Empirical $6.56+1.38$ $859.83+1.53$ $11.29+1.53$
Theoretical $0.24$ $0.94$ $0.30$
  • * Time of 2000 samples + inference.
Table 2 shows the average time of computing one $\rho(\varepsilon,\theta_{\diamond})$ and $\hat{\rho}(\varepsilon)$ for different mechanisms in this case study. For each empirical result $\hat{\rho}(\varepsilon)$, we sample $2000$ times from the mechanism and feed the samples to the classifier for inference. The theoretical approach is significantly faster, especially for the Exponential mechanism. For this mechanism, computing one empirical $\hat{\rho}(\varepsilon)$ requires $859.83$ milliseconds for $2000$ samples, whereas computing one theoretical $\rho(\varepsilon,\theta_{\diamond})$ takes only $0.94$ milliseconds, which is less than $0.1\%$. The efficiency stems from directly substituting the privacy parameter $\varepsilon$ and the robustness hyperrectangle $\theta_{\diamond}$ into the closed-form expression of $\rho(\varepsilon,\theta_{\diamond})$.
@@
7.2.2 Bank Customer Attrition

### v3

@@ 5.1. Robustness Hyperrectangle
The commonly used robustness radius for classifiers (DBLP:conf/nips/FawziMF16) is a conservative measure of robustness. It defines the robustness area as an $\ell_{\infty}$-ball, [^11] Other $\ell_{p}$-balls, such as $\ell_{1}$-ball and $\ell_{2}$-ball, are also used, but they are also symmetric across all dimensions. where all dimensions share the same radius. However, in practice, different dimensions often have varied robustness radii, making the true robustness area larger than that defined by an $\ell_{\infty}$-ball. In our context, accurately describing the robustness area is crucial for refining utility quantification. To this end, we define the robustness hyperrectangle, which accounts for the robustness of each dimension individually. The formal definition is as follows.
@@
The robustness radius is a special case of the robustness hyperrectangle, where each dimension has the same size $\theta$ around $x$. More generally, we say the classifier $h$ is $\theta_{\diamond}$-locally robust at $x$.
Finding robustness hyperrectangle. The robustness hyperrectangle $\theta_{\diamond}$ is not unique, unlike the robustness radius. Different methods for finding $\theta_{\diamond}$ may yield different results. Heuristically, we aim to include the robustness radius (i.e. $\theta\subseteq\theta_{\diamond}$) to ensure an improved utility quantification.
@@
Figure 5. Two robustness hyperrectangles (green dashed boxes) for the classifier in Figure 3. The original robustness radius $\theta=0.2$ corresponds to the gray dashed box.
@@
Each dimension of $\theta_{\diamond}$, represented as $a_{i},b_{i}]$, influences the concentration analysis of LDP mechanisms. Larger $a_{i},b_{i}]$ allows for greater perturbation, leading to a stronger utility guarantee. Specifically, the concentration level $\theta$ in the utility guarantee $\rho(\varepsilon,\theta)$ is replaced with the refined $\theta_{\diamond}$, resulting in a strictly improved utility guarantee $\rho(\varepsilon,\theta_{\diamond})$.
Refined utility quantification. Given a $d$-dimensional classifier $h$, LDP mechanism $\mathcal{M}_{\varepsilon}$ for each dimension, and sensitive data $x$. Denote the utility quantification $\rho(\varepsilon,\theta)$ as defined in Section 4.3. It can be refined to: With probability at least $\rho(\varepsilon,\theta_{\diamond})$, the classifier $h$ preserves the correct classification under the input perturbed by $d$ independent $\mathcal{M}_{\varepsilon}$ (i.e. pure $d\varepsilon$-LDP), where [^12] We omit the terms $(1-\omega)$ and $(1-\tau)$ related to $\rho(\varepsilon,\theta_{\diamond})$ for simplicity in presentation. They are almost $1$ in practice (e.g. $\omega=0.05,\tau=0.01$) and thus also negligible.
@@
5.2. PAC Privacy
Probably approximately correct (PAC) privacy (DBLP:conf/crypto/XiaoD23) provides a formal framework for relaxing the worst-case privacy guarantee. While pure LDP requires the privacy constraint to hold in all cases, PAC privacy allows a small probability of failure, i.e. probabilistic LDP. This relaxation allows the use of a wider range of mechanisms, notably the Gaussian mechanism, which is widely used in privacy-preserving machine learning.
To simplify notation, we denote the privacy loss (DBLP:journals/fttcs/DworkR14) (of a mechanism $\mathcal{M}$) between $x_{1}$ and $x_{2}$ at an observation $\tilde{x}$ as
@@ 5.2. PAC Privacy
Pure $\varepsilon$-LDP can then be expressed as: $\mathcal{L}_{\mathcal{M},x_{1},x_{2}}(\tilde{x})\leq\varepsilon$ for all $x_{1},x_{2},\tilde{x}$. In contrast, $(\varepsilon,\delta)$-PAC LDP is defined by a relaxed $\delta$-probabilistic condition. [^13] Readers may wonder about the relationship between $(\varepsilon,\delta)$-PAC LDP and $(\varepsilon,\delta)$-LDP. In fact, the traditional $(\varepsilon,\delta)$-LDP from Dwork (DBLP:journals/fttcs/DworkR14) is ambiguous and has many interpretations. Concentrated privacy (DBLP:journals/corr/DuchiWJ16) from Dwork, Rényi privacy (DBLP:conf/csfw/Mironov17), and PAC privacy (DBLP:conf/crypto/XiaoD23) can address this ambiguity. See Appendix C.4 for a detailed discussion.
Definition 0 (PAC LDP, adopted from (DBLP:conf/crypto/XiaoD23)).
@@
Combination of PAC LDP mechanisms. When applying $d$ independent PAC LDP mechanisms to $d$ dimensions, if we follow the combination theorem from Dwork (DBLP:journals/fttcs/DworkR14), it gives $(d\varepsilon,d\delta)$-PAC LDP. However, this result is not tight, as it allows the failure probability $d\delta$ to exceed $1$, which is impossible. We provide a tighter result for combining $d$ independent PAC LDP mechanisms.
@@ 5.2.1. Privacy Indicator
Refined utility quantification. Given a $d$-dimensional classifier $h$, combined LDP mechanism $\mathcal{M}^{d}_{\varepsilon}$, privacy indicator $\mathcal{I}_{\delta}(\cdot)$, and raw data $x$, the utility quantification $\rho(\varepsilon,\theta)$ can be refined to: With probability at least $\delta+(1-\delta)\rho(\varepsilon,\theta_{\diamond})$, the classifier $h$ preserves the correct classification under the input perturbed by $\mathcal{I}_{\delta}(\mathcal{M}^{d}(x))$, which guarantees $(d\varepsilon,\delta)$-PAC LDP.
@@
5.2.2. Extended Gaussian Mechanism
It is already known that the Gaussian mechanism (DBLP:journals/fttcs/DworkR14) cannot satisfy pure $\varepsilon$-LDP, making the privacy indicator inapplicable. Nonetheless, Dwork has shown that it can satisfy $(\varepsilon,\delta)$-LDP for $\varepsilon\leq 1$. Under the notion of PAC LDP, we provide an extended Gaussian mechanism to work for any $\varepsilon$. [^14] Analytical Gaussian mechanism (DBLP:conf/icml/BalleW18) works for any $\varepsilon$, but it is based on an alternative privacy definition and lacks analytical form of the noise scale $\sigma$. Appendix B.2.3 provides the detailed comparison. The following theorem is the form for data domain $x\in[0,1]$.
@@ 5.2.2. Extended Gaussian Mechanism
(Sketch) The proof generally follows the structure of Dwork et al (DBLP:journals/fttcs/DworkR14), but uses a rewritten privacy loss and a different tail bound for the Gaussian distribution. Appendix B.2 provides the details and compares this bound with others in the literature. ∎
@@
Refined utility quantification. Given a classifier $h$, combined Gaussian mechanism $\mathcal{M}^{d}$ with $\delta$, and $d$-dimensional sensitive data $x$, the utility quantification $\rho(\varepsilon,\theta_{\diamond})$ can be refined to: With probability at least $\rho(\varepsilon,\theta_{\diamond})$, the classifier $h$ preserves the correct classification under the input perturbed by $\mathcal{M}^{d}$, which guarantees $(d\varepsilon,1-(1-\delta)^{d})$-PAC LDP.
The extended Gaussian mechanism and the privacy indicator provide two approaches to achieve $(\varepsilon,\delta)$-PAC LDP. While both operate under the same privacy notion, the Gaussian mechanism actually provides a stronger effective privacy: when it fails to satisfy the pure $\varepsilon$-LDP (with probability $\delta$), the privacy loss exceeds $\varepsilon$ but remains bounded by a larger (unknown) value. In contrast, the privacy indicator exposes the original data when it fails.
6. Discussions
This section discusses detailed privacy questions in the utility quantification framework and summarizes other discussions that are deferred to the appendix.
@@ 6. Discussions
In the LDP paradigm, it is assumed that the classifier is curious about users’ sensitive data but honest in reporting results, i.e. it outputs $h(\mathcal{M}(x))$ rather than fabricating results. This is a weak assumption, as any utility evaluation becomes meaningless if the classifier fabricates results.
Utility quantification, as a different scenario from using the classifier, focuses on a specific input $x$ but does not involve interaction with users’ data. It can be conducted either by the classifier designer or by the users, and in both cases, there is no privacy leakage for the users. (i) If conducted by the classifier designer, e.g. for improving design, there is no interaction with users, and thus no privacy concerns arise. (ii) If conducted by the users, e.g. to achieve a better privacy-utility trade-off for their data, they can obtain the robustness radii without revealing their original data. For white-box classifiers, users know the classifier parameters and can compute the robustness radii. For black-box classifiers, users can uniformly query the classifier over the entire input domain (i.e. prediction for nearly all input $x\in[0,1]^{d}$), without revealing their specific data. This one-time procedure yields robustness radii for all inputs and privacy parameters. Thus, the robustness radius at any input can be considered public knowledge, which is a reasonable assumption.
Other Discussions
For brevity, additional discussions are deferred to Appendix C.5, including the following topics:
  1. (1)
    Complexity of finding the robustness radius.
  2. (2)
    Per-instance data utility vs aggregated data utility.
  3. (3)
    Fixed classifier with noisy input vs retrained classifier with noisy data.
  4. (4)
    Robustness of LDP mechanisms.
  5. (5)
    Independent vs correlated LDP mechanisms.
They provide additional insights into potential questions that may arise related to the utility quantification framework.
7. Case Studies
@@ 7.1. Setup
  • Stroke Prediction (stroke_dataset): Predicts stroke likelihood based on features like age, hypertension, and BMI. This dataset contains $6$ numerical input dimensions and a binary output. The sensitive features are “Age” and “BMI”.
  • Bank Customer Attrition (bank_attrition): Predicts whether a bank customer will leave based on features like age, salary, and tenure. This dataset has $9$ numerical input dimensions and a binary output. The sensitive features are “Age” and “Estimated Salary”.
  • MNIST-$\times7$ (variant of (DBLP:journals/spm/Deng12)): Predicts handwritten digits (0 to $9$) based on pixel values. This dataset has $49$ input dimensions and a $10$-class output. All dimensions are treated as sensitive features.
The first two datasets involve direct sensitive user inputs, while the third is a standard image classification dataset. For the first two datasets, we use the sklearn library with cross-validation to train Logistic Regression and Random Forest classifiers. These traditional classifiers are widely used in medical and financial domains due to their interpretability and lower data requirements (Amini25012023). For the third dataset, which consists of high-dimensional images, we use the torch library to train a Convolutional Neural Network (CNN) classifier. All datasets are normalized to $[0,1]$ for training, as normalization aids classifier convergence and is a common practice.
@@
Table 1. Summary of LDP mechanisms, types of classifiers, and datasets used in the case studies.
Mechanism Laplace (DBLP:journals/fttcs/DworkR14), Gaussian (Theorem 6), PM (DBLP:conf/icde/WangXYZHSS019), SW (DBLP:conf/sigmod/Li0LLS20), Exponential (DBLP:conf/innovations/NissimST12), $k$-$R$ (DBLP:conf/nips/KairouzOV14)
Classifier Low-dim: Logistic Regression, Random Forest; High-dim: Neural Network
Dataset Low-dim: Stroke Prediction (stroke_dataset), Bank Customer Attrition (bank_attrition); High-dim: MNIST-$\times7$ (variant of (DBLP:journals/spm/Deng12))
We present theoretical utility quantification for the classifiers under the LDP mechanisms discussed above. This quantification is based on the classifier’s robustness hyperrectangle $\theta_{\diamond}$ (at a record), determined using the search method described in Section 5.1. We also compare the theoretical results with empirical utility observations. Specifically, for a given classifier and mechanism, we provide and compare:
@@
7.2. Utility Quantification Results
We present the utility quantification results for the classifiers under input perturbation by different LDP mechanisms.
7.2.1. Stroke Prediction
@@
$\begin{aligned} \rho(\varepsilon,\theta_{\diamond})=&[F_{\text{PM}_{\varepsilon,\text{Age}}}(1)-F_{\text{PM}_{\varepsilon,\text{Age}}}(0.63)]\cdot[F_{\text{PM}_{\varepsilon,\text{BMI}}}(1)-F_{\text{PM}_{\varepsilon,\text{BMI}}}(0)]\\ =&1-F_{\text{PM}_{\varepsilon,\text{Age}}}(0.63).\end{aligned}$
@@
Figure 6. Empirical and theoretical utility for two classifiers trained on the Stroke Prediction dataset.
@@
The Random Forest classifier exhibits a larger gap between $\rho(\varepsilon,\theta_{\diamond})$ and $\hat{\rho}(\varepsilon)$ compared to the Logistic Regression classifier. This discrepancy arises from the robustness hyperrectangle $\theta_{\diamond}$, which is a conservative estimate of the overall robustness region (difficult to determine precisely). For Logistic Regression, the decision boundary is a hyperplane (sahu_decision_2024), making it more predictable and easier to approximate a hyperrectangle, resulting in a more accurate $\rho(\varepsilon,\theta_{\diamond})$. Appendix C.6.3 provides additional details on decision boundaries.
@@
Table 2 shows the average time of computing one $\rho(\varepsilon,\theta_{\diamond})$ and $\hat{\rho}(\varepsilon)$ for different mechanisms in this case study. For each empirical result $\hat{\rho}(\varepsilon)$, we sample $2000$ times from the mechanism and feed the samples to the classifier for inference. The theoretical approach is significantly faster, especially for the Exponential mechanism. For this mechanism, computing one empirical $\hat{\rho}(\varepsilon)$ requires $859.83$ milliseconds for $2000$ samples, whereas computing one theoretical $\rho(\varepsilon,\theta_{\diamond})$ takes only $0.94$ milliseconds, which is less than $0.1\%$. The efficiency stems from directly substituting the privacy parameter $\varepsilon$ and the robustness hyperrectangle $\theta_{\diamond}$ into the closed-form expression of $\rho(\varepsilon,\theta_{\diamond})$. Moreover, although computing $\rho(\varepsilon,\theta_{\diamond})$ requires the robustness hyperrectangle, $\theta_{\diamond}$ needs to be computed only once and can then be reused for all $\varepsilon$ values. In this case, computing $\theta_{\diamond}$ takes only $1.20$ milliseconds, which is still more efficient than estimating the empirical utility $\hat{\rho}(\varepsilon)$ separately for each $\varepsilon$.
Table 2. Time cost comparison (in milliseconds).
@@
Empiricala $6.56+1.38$ $859.83+1.53$ $11.29+1.53$
Theoreticalb $0.24$ $0.94$ $0.30$
  • a
    Time for generating 2000 samples + running inference.
  • b
    Time to compute $\rho(\varepsilon,\theta_{\diamond})$ only; computing $\theta_{\diamond}$ takes $1.20$ ms but is a one-time cost amortized across all $\varepsilon$ values.
@@
7.2.2. Bank Customer Attrition
Two key observations emerge in this case study: (i) The Exponential mechanism performs comparably to the PM mechanism for Logistic Regression. This is because its robustness hyperrectangle $\theta_{\diamond}=[0,0.72]\times[0,1]$ is large enough to cover most of the sensitive feature ranges, enabling excellent performance for all mechanisms defined on $[0,1]$. Figure 4 illustrates this phenomenon: larger $\theta$ values lead to $\rho(\varepsilon,\theta)$ approaching $1$ for all such mechanisms. However, Laplace and Gaussian mechanisms are less effective here due to their probability mass extending beyond $[0,1]$. (ii) The Gaussian mechanism is not as good as the other mechanisms. The main reason is from the noise scale $\sigma$ in Theorem 6. Although we can extend the Gaussian mechanism to $\varepsilon\geq 1$, this leads to a larger $\sigma$, i.e. less concentrated.